# Progressive Multimodal Search and Reasoning for Knowledge-Intensive Visual Question Answering

### v5

@@ 1 Introduction
We propose PMSR (Progressive Multimodal Search and Reasoning), which progressively constructs a structured reasoning trajectory to enhance knowledge acquisition and synthesis. Unlike prior approaches that condition each step on the full interaction history, PMSR maintains the reasoning state as a trajectory of compact records synthesized from retrieved evidence, and leverages this trajectory to guide subsequent retrieval and reasoning. To acquire diverse knowledge, PMSR formulates dual-scope queries conditioned on reasoning states at both the record and the trajectory levels. These queries then retrieve knowledge from heterogeneous knowledge bases (KBs), which PMSR synthesizes via compositional reasoning into a compact reasoning record. The updated record is appended to the reasoning trajectory to guide subsequent iterations. The process terminates adaptively when additional iterations become redundant.
@@ 2.1 Multimodal RAG
Multimodal retrieval-augmented generation (RAG) for knowledge-intensive VQA has largely followed a retrieve-then-read paradigm: external knowledge is retrieved in single-step and then read by the model to answer the question. Early work primarily focused on improving single-step retrieval by learning more effective multimodal embeddings Wei et al. (2024); Lin et al. (2024, 2025); Liu et al. (2024b); Jiang et al. (2025, 2024c). Subsequent approaches extended this paradigm by incorporating coarse-to-fine retrieval strategies. Hierarchical systems such as Wiki-LLaVA Caffagni et al. (2024) and EchoSight Yan and Xie (2024) adopt coarse-to-fine retrieval pipelines with image-based retrieval followed by multimodal or text-based reranking, while OMGM Yang et al. (2025) further develops this paradigm through a multi-step pipeline that explicitly models multiple knowledge granularities via successive multimodal and textual reranking.
@@ 2.2 Multimodal Agents
Building on these ideas, multimodal agents bring retrieval and tool use to vision–language settings by coupling retrieval with tool-augmented interaction, enabling models to iteratively reformulate queries and select external tools during multi-step reasoning. OmniSearch Li et al. (2024) introduces adaptive planning that routes multimodal queries across multiple search tools. More recent work learns search and tool-use policies via reinforcement learning: WebWatcher Geng et al. (2025) and MMSearch-R1 Wu et al. (2025) optimize retrieval trajectories over tool interactions, internalizing decision-making across steps. DeepEyesV2 Hong et al. (2025a) further integrates perception, search, and code execution within a unified agentic reasoning loop.
@@ 3.2 Dual-scope Query Formulation
After initialization, PMSR progressively guides knowledge search by generating new queries conditioned on the evolving reasoning trajectory. PMSR decomposes query generation into two complementary scopes: a record-level query grounded in the latest reasoning record and a trajectory-level query derived from the accumulated reasoning trajectory. The record-level query supports local refinement based on newly acquired knowledge, while the trajectory-level query preserves broader intent accumulated across iterations. The set of queries generated at iteration $t$ is given by:
@@
In contrast to the record-level query, it incorporates broader context accumulated over prior reasoning steps.
@@ 3.3 Joint Search across Heterogeneous KBs
where $\operatorname{sim}_{\text{img}}$ denotes cosine similarity in an image embedding space. We use a fixed weighting of $\lambda=0.5$. The text term adapts retrieval to the dual-scope query, while the image term preserves visual relevance to the input image.
@@ 4.1 Experiment Setup
Evaluation metrics.
@@
4.2 Retrieval Performance on VQA Benchmarks
[…]
As shown in Table 2, PMSR demonstrates strong and consistent retrieval performance on the OK-VQA benchmark, despite large differences from other benchmarks in knowledge type, question formulation, and grounding requirements. On OK-VQA, PMSR achieves 92.1% and 97.1% cumulative recall with the 4B and 8B models, closely matching its performance on InfoSeek and E-VQA. This cross-domain stability contrasts sharply with prior retrieval-augmented models, which often perform well only within their target domain. The results indicate that PMSR’s progressive retrieval strategy generalizes effectively across knowledge types without requiring dataset-specific tuning.
@@ 4.3 Accuracy on Knowledge-Intensive VQA
4.4 Results on Search-Oriented Benchmarks
Method FVQA test InfoSeek Human MM Search Live VQA
OmniSearch (GPT-4o) Li et al. (2024) - - 49.7 40.9
MMSearch-R1 Wu et al. (2025) 58.4 55.1 53.8 48.4
WebWatcher Geng et al. (2025) - - 49.1 51.2
DeepEyesV2 Hong et al. (2025a) 60.6 51.1 63.7 -
Ours 61.2 58.2 54.3 54.2
Table 4: Performance on search-oriented multimodal benchmarks. Results for OmniSearch are taken from WebWatcher Geng et al. (2025). Unless otherwise noted, all methods use Qwen2.5-VL-7B as the backbone.
We further evaluate PMSR on search-oriented benchmarks that require multimodal grounding and open-domain knowledge acquisition. As shown in Table 4, using the same Qwen2.5-VL-7B backbone, PMSR achieves 61.2% and 58.2% accuracy on FVQA and InfoSeek Human, respectively, surpassing recent agent-based baselines. On MMSearch, PMSR attains 54.3% accuracy, remaining competitive with specialized multimodal search agents. On LiveVQA, which targets real-world, time-sensitive information seeking over diverse news sources, PMSR reaches 54.2% accuracy, the highest among the methods reported in Table 4.
@@ 5 Ablations
For ablation studies, we conduct experiments using Qwen3-VL-8B to validate the robustness of individual components. Unless otherwise specified, the retrieval budget is fixed across single-query and dual-scope query settings, and adaptive termination is used with $\tau=0.9$ (up to a maximum of 5 iterations).
@@ 8 Limitations
While the PMSR framework demonstrates significant improvements in retrieval recall and answer accuracy across several knowledge-intensive VQA benchmarks, several limitations warrant consideration for future research. First, the proposed framework relies on iterative retrieval and reasoning, which introduces additional inference overhead compared to single-pass RAG methods. Although the adaptive termination mechanism mitigates redundant iterations, the overall computational cost remains higher in cases where convergence is slow.
@@ Appendix I Record-level Query Analysis
Appendix P Computational Cost Analysis
We analyze the per-sample computational cost of PMSR with three iterations on a commodity server (NVIDIA RTX 3090, Intel Xeon Gold 6248R) using local retrieval. Under this setup, PMSR requires 17.2 seconds per iteration to process one sample end-to-end. Importantly, this latency is not inherent to the framework but mainly reflects an unoptimized implementation.
The overhead stems primarily from (i) serial query execution across multiple knowledge bases with vanilla index configurations, and (ii) model inference without an optimized serving stack (e.g., vLLM or SGLang). Both factors are amenable to system-level optimization, such as parallelizing retrieval across sources, tuning/accelerating indexing and search, and using optimized inference runtimes.

### v6

@@ 1 Introduction
We propose PMSR (Progressive Multimodal Search and Reasoning), which progressively constructs a structured reasoning trajectory to enhance knowledge acquisition and synthesis. Unlike prior approaches that condition each step on the full interaction history, PMSR maintains the reasoning state as a trajectory of compact records synthesized from retrieved evidence, and leverages this trajectory to guide subsequent retrieval and reasoning. Specifically, PMSR is built on two key ideas: record-isolated updates, where each iteration synthesizes a new reasoning record solely from newly retrieved evidence, and dual-scope querying, which decouples the latest reasoning state from the overall trajectory to support both local retrieval refinement and trajectory-level reflection. To acquire diverse knowledge, PMSR formulates dual-scope queries to retrieve complementary evidence from heterogeneous knowledge bases (KBs). The retrieved evidence from diverse sources is synthesized through compositional reasoning into a compact reasoning record and appended to the trajectory for the next iteration.
@@ 2.1 Multimodal RAG
Multimodal retrieval-augmented generation (RAG) for knowledge-intensive VQA has largely followed a retrieve-then-read paradigm: external knowledge is retrieved in single-step and then read by the model to answer the question. Early work primarily focused on improving single-step retrieval by learning more effective multimodal embeddings Wei et al. (2024); Lin et al. (2024, 2025); Liu et al. (2024b); Jiang et al. (2025b, 2024c). Subsequent approaches extended this paradigm by incorporating coarse-to-fine retrieval strategies. Hierarchical systems such as Wiki-LLaVA Caffagni et al. (2024) and EchoSight Yan and Xie (2024) adopt coarse-to-fine retrieval pipelines with image-based retrieval followed by multimodal or text-based reranking, while OMGM Yang et al. (2025) further develops this paradigm through a multi-step pipeline that explicitly models multiple knowledge granularities via successive multimodal and textual reranking.
@@ 2.2 Multimodal Agents
Building on these ideas, multimodal agents bring retrieval and tool use to vision–language settings by coupling retrieval with tool-augmented interaction, enabling models to iteratively reformulate queries and select external tools during multi-step reasoning. OmniSearch Li et al. (2024) introduces adaptive planning that routes multimodal queries across multiple search tools. More recent work learns search and tool-use policies via reinforcement learning: WebWatcher Geng et al. (2025) and MMSearch-R1 Wu et al. (2025) optimize retrieval trajectories over tool interactions, internalizing decision-making across steps. DeepEyesV2 Hong et al. (2025a) further integrates perception, search, and code execution within agentic loop.
@@ 3.2 Dual-scope Query Formulation
After initialization, PMSR progressively guides knowledge search by generating new queries conditioned on the evolving reasoning trajectory. PMSR decomposes query generation into two complementary scopes: a record-level query grounded in the latest reasoning record and a trajectory-level query derived from the accumulated reasoning trajectory. The record-level query supports local refinement by using the latest reasoning record to retrieve evidence closely related to the most recent deduction, whereas the trajectory-level query supports global reflection by analyzing compact records in the trajectory to identify unresolved gaps, resolve conflicts, and retrieve broader contextual evidence.
The set of queries generated at iteration $t$ is given by:
@@
In contrast to the record-level query, it incorporates broader context accumulated over prior reasoning steps, where its goal is to guide the proper direction for the next query given reasoning records.
@@ 3.3 Joint Search across Heterogeneous KBs
where $\operatorname{sim}_{\text{img}}$ denotes cosine similarity in an image embedding space, and we use a fixed weight of $\lambda=0.5$ to balance the two modalities. The text term adapts retrieval to the dual-scope query, while the image term preserves visual relevance to the input image.
@@ 4.1 Experiment Setup
4.2 Retrieval Performance on VQA Benchmarks
[…]
As shown in Table 2, PMSR demonstrates strong and consistent retrieval performance on the OK-VQA benchmark, despite highly different from other benchmarks in knowledge type, question formulation, and grounding requirements. On OK-VQA, PMSR achieves 92.1% and 97.1% cumulative recall with the 4B and 8B models, closely matching its performance on InfoSeek and E-VQA. This cross-domain stability contrasts sharply with prior retrieval-augmented models, which often perform well only within their target domain. The results indicate that PMSR’s progressive retrieval strategy generalizes effectively across knowledge types without requiring dataset-specific tuning.
4.3 Accuracy on Search Benchmarks
Method FVQA test InfoSeek Human MM Search Live VQA
OmniSearch (GPT-4o) Li et al. (2024) - - 49.7 40.9
MMSearch-R1 Wu et al. (2025) 58.4 55.1 53.8 48.4
WebWatcher Geng et al. (2025) - - 49.1 51.2
DeepEyesV2 Hong et al. (2025a) 60.6 51.1 63.7 -
Ours 61.2 58.2 54.3 54.2
Table 3: Performance on search-oriented multimodal benchmarks. Results for OmniSearch are taken from WebWatcher Geng et al. (2025). Unless otherwise noted, all methods use Qwen2.5-VL-7B as the backbone, ensuring a fair comparison.
We evaluate PMSR on search-oriented benchmarks that require multimodal grounding and open-domain knowledge acquisition. As shown in Table 3, using the same Qwen2.5-VL-7B backbone, PMSR achieves 61.2% and 58.2% accuracy on FVQA and InfoSeek Human, respectively, surpassing recent agent-based baselines. On MMSearch, PMSR attains 54.3% accuracy, remaining competitive with specialized multimodal search agents. On LiveVQA, which targets real-world, time-sensitive information seeking over diverse news sources, PMSR reaches 54.2% accuracy, the highest among the methods reported in Table 3.
@@ 5 Ablations
For ablation studies, we conduct experiments using Qwen3-VL-8B to validate the robustness of individual components. Unless otherwise specified, the retrieval budget is fixed across single-query and dual-scope query settings, and adaptive termination is used with $\tau=0.9$ (up to a maximum of 5 iterations). Specifically, under the heterogeneous KB setting, we retrieve a total of 20 text passages and 10 image-text pairs per iteration; for dual-scope querying, this budget is evenly split across the two queries. To maintain a comparable computational budget, when using only the multimodal KB, we retrieve 20 image-text pairs per iteration.
@@ 8 Limitations
While the PMSR framework demonstrates significant improvements in retrieval recall and answer accuracy across several knowledge-intensive VQA benchmarks, few limitations warrant consideration for future research. First, the proposed framework relies on iterative retrieval and reasoning, which introduces additional inference overhead compared to single-pass RAG methods. Although the adaptive termination mechanism mitigates redundant iterations, the overall computational cost remains higher in cases where convergence is slow.
@@
Acknowledgments
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) ([NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)], [No.RS-2023-00235293, Development of autonomous driving big data processing, management, search, and sharing interface technology to provide autonomous driving data according to the purpose of usage]).
@@ Appendix I Record-level Query Analysis
Method Stable-Correct Correction Conflicts Persistent-Fail
PMSR (Ours) 47.7 30.1 17.3 4.9
WebWatcher 12.1 31.2 4.9 51.8
Table A9: Trajectory-type distribution on E-VQA test split. Trajectory types are defined by per-iteration correctness of records(BEM): Stable-Correct (correct at all iterations), Persistent-Fail (incorrect at all iterations), Correction (recovers from incorrect reasoning and is correct at the final iteration), Conflicts (correct in some iterations but incorrect at the final iteration).
@@
Appendix P Runtime Latency Analysis
To quantify the overhead of progressive search and reasoning, we report wall-clock latency under the same inference setting used in our main experiments. Unless otherwise stated, all times are reported in seconds. We first measure a naïve sequential implementation on a commodity server with a single RTX 3090, and then evaluate two practical optimizations: concurrent retrieval across heterogeneous KBs and faster VLM decoding on a higher-capability GPU.
P.1 Runtime Latency Breakdown
Operation Type Component Time (s)
Reasoning Dual-Scope Query Formulation 3.34
Retrieval Two Text Index Retrievals 5.51
Retrieval Two Multimodal Index Retrievals 2.83
Reasoning Reasoning Record Generation 5.52
Table A10: Component-wise latency breakdown of PMSR for one reasoning iteration under a vanilla implementation.
Setting RAG (1-step) PMSR (1-iter.)
Baseline 3.91 17.20
+ Concurrent retrieval 3.91 11.61
+ H100 GPU decoding 2.14 4.33
Table A11: Runtime comparison between single-step RAG and PMSR. All times are reported in seconds.
Under the naïve sequential setting, the per-iteration runtime is 17.20 seconds. Table A10 provides a component-wise latency analysis of PMSR for one reasoning iteration, showing that retrieval and VLM decoding account for most of the latency.
P.2 Mitigating Runtime Latency
Table A11 summarizes the effect of two practical runtime optimizations for PMSR. First, textual and multimodal retrieval can be executed concurrently, since PMSR uses dual-scope queries to retrieve knowledge over heterogeneous sources. This reduces retrieval wall-clock time from 8.34 seconds to 2.75 seconds and lowers the per-iteration latency of PMSR from 17.20s to 11.61s without changing the algorithm itself.
Second, VLM decoding latency can be accelerated by using a higher-capability GPU (Nvidia H100). Using a higher-capability GPU for VLM decoding reduces the latency of Qwen3-VL-8B from 8.86 seconds to 1.58 seconds per sample, further lowering PMSR latency to 4.33 seconds per iteration. Overall, these practical optimizations reduce PMSR latency as 4.33 seconds per iteration, substantially narrowing the gap to single-step RAG. With adaptive termination, PMSR executes 3.5 iterations on average, corresponding to an end-to-end runtime of approximately 15.15 seconds per sample.
Appendix Q Comparison of Record-level Query Design
Record-level query operator Dataset Acc. ($\uparrow$) Recall ($\uparrow$)
Concatenation InfoSeek 53.5 95.5
Entity extraction InfoSeek 54.5 94.9
Concatenation E-VQA 43.5 66.7
Entity extraction E-VQA 41.0 64.9
Table A12: Comparison of record-level query operators. Results are reported on subsets of the InfoSeek validation set and the E-VQA test split for efficiency.
We also analyze the design of the record-level query transformation. In PMSR, the record-level query is formed by concatenating the original question with the latest reasoning record. To evaluate whether a complex transformation is beneficial, we additionally compare concatenation with an entity-extraction-based operator on InfoSeek and E-VQA. As shown in Table A12, the two operators yield comparable results on InfoSeek, while concatenation performs better on E-VQA. These results suggest that a more complex query transformation is not consistently beneficial in our setting. Since the reasoning record already captures the latest deductions from compositional reasoning, and converting it into an alternative query form remains challenging.
Appendix R Analysis of Compositional Reasoning in Reasoning Record Generation
Dataset Filter Summarize Compose Resolve Conflicts
InfoSeek 19.0 92.0 48.0 62.0
E-VQA 23.5 86.0 50.5 70.5
Table A13: LLM-as-a-judge analysis of behaviors exhibited by reasoning records generated by $G_{\text{reason}}$. Each trajectory can be assigned multiple labels, since its reasoning records can jointly filtering, summarization, composition, and resolving conflicts across iterations.
To assess whether the $G_{\text{reason}}$ operator performs compositional reasoning, we conduct an LLM-as-a-judge analysis over reasoning records generated by PMSR trajectories. For each sample, the judge assigns one or more behavior labels from Filter, Summarize, Compose, and Resolve Conflicts. Here, Compose denotes combining multiple pieces of evidence into a conclusion not explicitly stated in any single retrieved item, while Resolve Conflicts denotes identifying or reconciling inconsistent evidence across retrieved candidates.
As shown in Table A13, Summarize is common, as the reasoning record consolidates retrieved knowledge into a compact state. At the same time, we observe substantial rates of Compose and Resolve Conflicts: 48.0% and 62.0% on InfoSeek, and 50.5% and 70.5% on E-VQA, respectively. These results indicate that $G_{\text{reason}}$ performs compositional reasoning to synthesize information beyond summarization.
Appendix S Comparison with MLLM-based Retrievers
Retriever Dataset Acc. ($\uparrow$) Recall ($\uparrow$)
Default retriever (ours) InfoSeek 53.5 95.5
Qwen3-VL-Embedding-2B InfoSeek 53.1 92.9
Default retriever (ours) E-VQA 43.5 66.7
Qwen3-VL-Embedding-2B E-VQA 43.3 60.0
Table A14: Comparison between the default retriever used in PMSR and an MLLM-based retriever. Results are reported on subsets of the InfoSeek validation set and the E-VQA test split for efficiency.
Our main experiments use standard dense retrievers for both textual and multimodal retrieval. We adopt this setting because PMSR requires pairwise retrieval over multimodal KBs, which many recent MLLM-based retrievers do not explicitly target. However, some recent MLLM-based retrievers support the pairwise retrieval capability required by PMSR. To examine whether such retrievers provide complementary gains in our framework, we additionally integrate Qwen3-VL-Embedding-2B Li et al. (2026) into PMSR and evaluate it on InfoSeek and E-VQA.
As shown in Table A14, the MLLM-based retriever does not consistently improve either retrieval recall or end-to-end answer accuracy over the default retriever. On InfoSeek, Qwen3-VL-Embedding-2B achieves 53.1 accuracy and 92.9 recall, compared to 53.5 and 95.5 for the default retriever; on E-VQA, it achieves 43.3 accuracy and 60.0 recall, compared to 43.5 and 66.7. These results show that, under our current setup, our retriever configuration yields better results than the similarly sized MLLM-based retriever.
Appendix T Generalizability Beyond Knowledge-Intensive VQA
Method Acc. ($\uparrow$)
CoT 31.1
RAG 46.6
IRCoT Trivedi et al. (2023) 50.9
Search-R1-7B Jin et al. (2025) 58.6
s3 Jiang et al. (2025a) 59.0
PMSR (ours) 59.8
Table A15: Cross-domain evaluation on the HotpotQA dev set under an LLM-based evaluation protocol, following the LLM-as-a-judge protocol used in s3.
To examine the cross-domain applicability of PMSR beyond multimodal VQA, we additionally evaluate PMSR on text-only knowledge-intensive question answering. Specifically, we use the HotpotQA Yang et al. (2018) dev split as a representative benchmark and adapt PMSR to the text-only setting while preserving its core design: progressive retrieval, record-level and trajectory-level query formulation, and structured reasoning-state updates over a textual KB only.
As shown in Table A15, PMSR achieves competitive performance on HotpotQA and outperforms several iterative retrieval baselines. These results suggest that the proposed framework is applicable beyond visual question answering and can also be effective in broader knowledge-intensive domains.